The curve of a 70.0-m circular path is banked at 35O. At what velocity shall a 1000- kg car have for it to make the turn in this curve? 




1
Expert's answer
2021-12-01T10:35:58-0500

reaction of road in x-direction:

"N=mgcos10\degree sin10\degree"


then:

"N=mv^2\/R"


"v=\sqrt{RN/m}=\sqrt{gRcos10\degree sin10\degree}=\sqrt{70gcos10\degree sin10\degree}=10.83" m/s